OpenStudy (anonymous):

What is the value of n? if 3^10*27^2=9^2*3^n

OpenStudy (uri):

just got an headache,a really bad one.

OpenStudy (lncognlto):

Where is n in the question?

OpenStudy (uri):

^

OpenStudy (anonymous):

its 3^n

OpenStudy (uri):

Oh.

OpenStudy (lncognlto):

Sorry, right, here goes

OpenStudy (lncognlto):

3^7290 = 9^2*3^n 7290 log 3 = 2*3^n (log 9) (all base 10) 3645 log 3 = 3^n log 9 3^n = (3645 log 3/log 9) n log 3 = log (3645 log 3/log 9) n = (log (3645 log 3/log 9))/log 3 There we go, that should do it. Now we evaluate with a calculator. The principle I am using is that x^n = n log x.

OpenStudy (anonymous):

Is that the correct method for solving these kinda questions?

OpenStudy (anonymous):

its werid......Easy way is to make the powers on both side same and then equate powers

OpenStudy (anonymous):

Can you solve? this question..Then I would understand how to solve these questions!!

OpenStudy (anonymous):

\[ 3^(10)*27^2=9^2*3^n\] 3^10 * 3^6 = 3^4 * 3^n 16 =4+n n=12

OpenStudy (anonymous):

27 = 3^3 9 = 3^2

OpenStudy (anonymous):

Got it! Thanks..

OpenStudy (anonymous):

Welcome..:)

OpenStudy (debbieg):

Yeah, was just coming to post what @Yahoo! did. It's easier than you think, just convert all to base 3 and combine and you end up with something of the form\[\Large 3^a=3^b\] Hence, a=b and it's an easy-peasy linear equation. :)